A firm that is in the 35% tax bracket forecasts that it can retain $4 million of new earnings plans to raise new capital in the following proportions:

60% from 30-year bonds with a flotation cost of 4% of face value. Their current bonds are selling at a price of 91 (91% of face value), have 4 years remaining, have an annual coupon of 7%, and their investment bank thinks that new bonds will have a 40 basis point (0.40%) higher yield-to-maturity than their current 4-year bonds due to their longer term. Any new bonds will be sold at par.
10% from preferred stock with a flotation cost of 5% of face value. The firm currently has an outstanding issue of $30 face value fixed-rate preferred stock with an annual dividend of $2 per share, and the stock is currently selling at $27 per share. Any newly issued preferred stock will continue with the $30 par-value, and will continue with the $2 dividend.
30% from equity. Their common dividend payout ratio is 60%, they paid a dividend of $1.59 per share yesterday, the dividend is expected to grow to $4.22 in 20 years, and is expected to continue this growth rate into the foreseeable future. The common stock has a current market price of $19, and their investment banker suggests a flotation cost of 7% of market value on new common equity.


Part 1: Calculate the after-tax cost of the new bond financing. ___________

Part 2: Calculate the after-tax cost of the new preferred stock financing. ______

Part 3: Calculate the after-tax cost of retained earnings financing. _______

Part 4: Calculate the after-tax cost of the new common equity financing. ______
Answer #1

NIMER: 2.7 POLO TË Current Price = = 914 frotation costi 4 =95 S t Coupon i piments Maturing price n = 100. ATM CTML MAP 2 1

Page Date Porfeded Stoon Cost Finuncing divident = 251 + Mayket vanie = 27 +2221$ Cost of I - Poefedence_ Divo 100 Ro 2 x bo

(d) cost of Eauity the = Dividenta Po 1.67 x 100+5 20.33 = 18,21% 7+5%. = 113.21.17 (2 cost of Retained Earninys: .- Retainedthank you for posting

I hope my efforts will be fruitful to you...?